Po²¹² undergoes alpha decay. Alpha particle have K.E of 8.776MeV. Assume the mass of Alpha particle to be 6.67*10^(-27). Calculate their velocity
1
Expert's answer
2020-04-20T19:04:12-0400

KE = "\frac12mv^2"

"8.776\times10^6\times1.6\times10^{-19}=\frac12\times6.67\times10^{-27}\times v^2"

"4.2\times10^{14}=v^2\\v=2.05\times10^7m/s"